OpenStudy (anonymous):

The sum of two numbers is 57. One of the numbers is 21. Which of the following equations could be used to find the other number?

OpenStudy (anonymous):

\[\sum = number_1 + number_2\]\[57 = 21 + number_2\]\[number_2 = 57 - 21 = 36\] of course, this could also be "number 1" ;)

OpenStudy (anonymous):

what are the choices

OpenStudy (anonymous):

57 = n + 21 57 = 2 + 21 21 = n – 57 n = 57 + 21 Not Here

OpenStudy (anonymous):

carolsan: the first equation is the correct one. :)

OpenStudy (anonymous):

the first

OpenStudy (anonymous):

cool thx guys